# DynamicRAG: Leveraging Outputs of Large Language Model as Feedback for Dynamic Reranking in Retrieval-Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) systems combine large language models (LLMs) with external knowledge retrieval, making them highly effective for knowledge-intensive tasks. A crucial but often under-explored component of these systems is the reranker, which refines retrieved documents to enhance generation quality and explainability. The challenge of selecting the optimal number of documents (k) remains unsolved: too few may omit critical information, while too many introduce noise and inefficiencies. Although recent studies have explored LLM-based rerankers, they primarily leverage internal model knowledge and overlook the rich supervisory signals that LLMs can provide, such as using response quality as feedback for optimizing reranking decisions. In this paper, we propose DynamicRAG, a novel RAG framework where the reranker dynamically adjusts both the order and number of retrieved documents based on the query. We model the reranker as an agent optimized through reinforcement learning (RL), using rewards derived from LLM output quality. Across seven knowledge-intensive datasets, DynamicRAG demonstrates superior performance, achieving state-of-the-art results. The model, data and code are available at https://github.com/GasolSun36/DynamicRAG.
@@ 1 Introduction
Retrieval-augmented generation (RAG) systems have emerged as a powerful approach for combining the strengths of large language models (LLMs) with external knowledge retrieval. This integration has proven highly effective for addressing knowledge-intensive tasks and incorporating up-to-date information into LLMs, leading to notable performance improvements Izacard et al. (2023); Kulkarni et al. (2024); Guu et al. (2020). Consequently, RAG systems have garnered considerable interest from both the academic and industrial communities. A crucial, yet often underappreciated, component of RAG systems is the reranker, which assesses the relevance of retrieved documents. The reranker is critical for improving the quality of generated text and enhancing explainability, thereby serving as an indispensable part of the RAG framework Ma et al. (2023); Pradeep et al. (2023).
@@
In RAG systems, the reranker’s primary role is to refine the Top-N documents retrieved by the retriever, selecting the k most relevant ones to enhance the answer quality. However, determining the optimal k remains a challenging problem, as highlighted in previous studies Nguyen et al. (2024); Yu et al. (2024b). A k that is too small risks omitting critical information, leading to degraded generation quality, while a larger k may introduce irrelevant content, increasing noise and potentially misleading the generator. Furthermore, incorporating excessively long contexts can reduce both efficiency and effectiveness, further complicating the balance between relevance and performance in RAG systems. Striking the right balance requires adaptive strategies that can dynamically adjust k based on query complexity and document diversity.
@@
Figure 1: Illustration of our dynamic reranker framework. Figure (a) represents a RAG system without a reranker, where the system primarily focuses on training LLMs. Figure (b) represents a RAG system with a reranker, where the reranker is trained to filter the Top-N documents to a fixed Top-K, which remains constant for all queries. In contrast, Figure (c) represents our dynamic reranker, where both the reranker and the generator are trained simultaneously. The dynamic reranker adapts to the difficulty of each query by dynamically determining the value of k.
Recent studies Drozdov et al. (2023); Sun et al. (2023); Lewis et al. (2020) have explored the use of LLMs as rerankers, leveraging their ability to generate rankings of documents. For example, Khramtsova et al. (2024) proposes using LLMs to generate synthetic queries for the target corpus to improve the ranking of documents. Liu et al. (2024a) introduces DemoRank, a two-stage "retrieve-then-rerank" framework, where LLMs score demonstrations and iteratively refine the ranking. Other works have also used the understanding capabilities of LLMs through sliding window mechanisms to achieve optimal re-ranking results Sun et al. (2023); Chao et al. (2024). These studies show the versatility and potential of LLMs in improving the reranking task in different contexts. However, previous studies primarily exploit the internal knowledge of LLMs but overlook the rich supervisory signals that LLMs can provide. For instance, the quality of an LLM’s response—given a query and a ranked set of documents—serves as a direct indicator of document relevance.
@@
We comprehensively evaluate DynamicRAG on knowledge-intensive tasks using seven datasets: NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), HotpotQA Yang et al. (2018), 2WikimQA Ho et al. (2020), ASQA Stelmakh et al. (2022), FEVER Thorne et al. (2018), and ELI5 Fan et al. (2019). Additionally, we assess the recall results of DynamicRAG’s reranker on NQ and HotpotQA. Experimental results show that DynamicRAG significantly outperforms existing fine-tuned and prompting-based approaches, achieving state-of-the-art (SOTA) performance.
@@ 2.1.2 Encoding Phase
Once the relevant documents have been retrieved, both the query $\mathbf{q}$ and the documents $\{D_{1},D_{2},\dots,D_{k}\}$ are encoded into fixed-size vectors using a neural encoder such as BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), or other transformer-based models. The goal is to obtain a dense representation of both the query and documents. Let $\mathbf{q}_{\text{enc}}$ and $\mathbf{D}_{i}$ represent the encoded query and the encoding of document $D_{i}$, respectively:
@@ 3 DynamicRAG
The DynamicRAG framework consists of three key components: (1) a frozen retriever, (2) a trainable dynamic reranker, and (3) a trainable generator. The retriever retrieves relevant documents from a large corpus, while the reranker dynamically determines both the order and the number of documents to be passed to the generator to produce an answer. The generator then produces the final output based on the reranker’s selected documents. By iteratively training the reranker and generator, DynamicRAG achieves improvements in the overall efficiency and effectiveness of the RAG system.
@@ 3.1 Dynamic Reranking
Traditional reranking approaches rely on static ranking models that determine the relevance of retrieved documents independently of the generation task. These models typically operate with a fixed length, where the input consists of $n$ documents, and the output is a reordered sequence of the same $n$ documents. In contrast, DynamicRAG uses feedback from the generator to guide the reranking process, allowing it to dynamically adjust both the order and the number of documents selected. This enables the reranker to optimize the input of the generator, maximizing the likelihood of producing high-quality output [^1] In this paper, we only consider the list-wise ranking..
To formalize, the reranking process can be expressed as a single operation that directly outputs a reordered subset of the retrieved documents. Given a query $\mathbf{q}$ and a set of retrieved documents $D=\{D_{1},D_{2},\dots,D_{k}\}$, the reranker computes:
@@
where $\hat{D}\subset D$ is the selected and reordered subset of documents, and $\theta_{r}$ represents the parameters of the reranker. This formulation encapsulates both the scoring and selection processes, producing a subset $\hat{D}$ with dynamically adjusted order and size.
@@ 3.2 Training with Interacting
In this section, we outline the training process for the Reranker. The training begins with behavior cloning, which equips the Reranker with the basic ability to adjust both the order and the number of selected documents. Building upon this foundational model, we further refine the Reranker’s performance through interactions with the environment. This interaction allows the Reranker to collect feedback from multiple trajectories and enhance its decision-making policy. The complete training procedure is detailed in Algorithm 1.
@@ 3.2.1 Behavioral Cloning with Expert Trajectories
Behavioral cloning is used to supervised fine-tuning the Reranker by mimicking expert trajectories, allowing the model to learn the fundamental actions required for effective ranking. In this stage, the Reranker focuses on learning how to predict the correct intermediate actions $a_{t}$ based on the given task and context.
To achieve this, the Reranker is trained on a dataset of expert demonstrations, denoted as $D_{e}$. This enables the model to acquire basic instruction-following capabilities and leverage prior knowledge. The training objective is to maximize the likelihood of the expert actions at each time step:
$\mathcal{J}_{BC}(\theta)=\mathbb{E}_{(\mathcal{G},i,\tau)\sim\mathcal{D}_{e}}\Bigg{[}\sum_{t=1}^{T}\log\pi_{\theta}(a_{t}|\mathcal{G},i,h_{t-1})\Bigg{]},$ (10)
where $\mathcal{G}$ denotes the environment, $i$ denotes the user instruction, $h_{t-1}$ denotes the accumulated context until time step $t-1$, and $\pi_{\theta}(a_{t}|\mathcal{G},i,h_{t-1})$ is the conditional probability of taking action $a_{t}$ given the current state.
@@ 3.2.2 Optimizing the Reranker via Exploration and Feedback
where $\mathbb{R}(\mathcal{G},i,\tau)$ quantifies the response quality. However, optimizing this objective is challenging due to the complexity of the action space and the inherent inefficiencies in sampling. To address this, we adopt the Direct Preference Optimization (DPO) framework Rafailov et al. (2024), which simplifies the reward optimization with pairwise comparisons of trajectories.
@@ 3.2.3 Reward Function Design
  • Semantic Similarity (SS): Assesses the alignment between $\hat{y}$ and $y_{gt}$ using BERTScore Zhang et al. (2020):
@@
  • Textual Fluency (TF): Evaluates fluency using ROUGE Lin (2004) metrics:
@@
  • LLM-Based Evaluation (LLM-Eval): Uses LLM-based scoring to assess alignment with task requirements, where $\mathcal{P}$ denotes the scoring prompt, detailed in Appendix B.6:
@@
3.3 Generation with Reranked Contexts
The generator produces high-quality responses using reranked documents $\hat{D}$. It is trained to maximize the likelihood of generating accurate outputs $\hat{y}$ aligned with ground-truth answers $y_{gt}$:
$\mathcal{J}_{\text{gen}}(\theta_{g})=\mathbb{E}_{(\mathbf{q},\hat{D},y_{gt})\sim\mathcal{D}_{\text{train}}}\Bigg{[}\sum_{t=1}^{T}\log p_{\theta_{g}}(y_{t}\mid\mathbf{q},\hat{D},y_{<t})\Bigg{]},$ (21)
where $\theta_{g}$ represents generator parameters, and $y_{t}$ denotes the token at time step $t$.
@@ 3.3 Generation with Reranked Contexts
IRCoT Trivedi et al. (2023) N/A - - 17.4 - - - -
ReAct Yao et al. (2023) N/A - - 35.1 - - 62.0 -
RA-DIT Lin et al. (2024) $\sim$ 1,129k 43.5 72.8/- 36.6 - - 86.9 -
FLARE* Jiang et al. (2023) N/A - - - 51.0 41.3 - -
Reward-RAG Nguyen et al. (2024) Unknown 42.2 75.6/80.4 - - - 89.8 -
LLaMA2-7B Touvron et al. (2023) N/A 17.9 -/42.5 16.6 17.9 19.0 30.0 15.6
@@
Self-RAG (LLaMA2-7B) Asai et al. (2023) $\sim$ 150k 36.4 -/66.4 - - 30.0 - -
DRAGIN Su et al. (2024) N/A - - 23.2 22.0 - - -
Smart-RAG (LLaMA2-7B) Gao et al. (2024) $\sim$ 218k - - - - 26.6 - -
@@
Auto-RAG (LLaMA3-8B-Instruct) Yu et al. (2024a) Unknown 37.9 60.9/- - - 30.0 - -
ChatQA-1.5 (LLaMA3-8B) Liu et al. (2024b) $\sim$ 442k 42.4 81.0/87.6 33.4 26.8 - 90.9 -
RankRAG (LLaMA3-8B) Yu et al. (2024b) $\sim$ 470k 50.6 82.9/89.5 35.3 31.4 - 93.8 -
@@
3.4 Inference Phase
During inference, the retriever selects documents $D$, and the reranker refines them into an ordered subset $\hat{D}$:
$\hat{D}=\texttt{Reranker}_{\theta_{r}}(\mathbf{q},D).$ (22)
The generator then produces the response:
$\hat{y}=\text{argmax}_{y}\,p(y\mid\mathbf{q},\hat{D}).$ (23)
This pipeline ensures the generation of high-quality, contextually relevant answers.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate DynamicRAG’s performance on comprehensive knowledge-intensive question-answering tasks, spanning over seven datasets and covering different types of challenges: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), HotpotQA Yang et al. (2018), 2WikimQA Ho et al. (2020), FEVER Thorne et al. (2018), ASQA Stelmakh et al. (2022), and ELI5 Fan et al. (2019). For the first five datasets (NQ, TriviaQA, HotpotQA, 2WikimQA and ASQA), we follow prior studies and adopt exact match (Hits@1) as the evaluation metric Asai et al. (2023); Yu et al. (2024b). For TriviaQA and FEVER, we used accuracy, while for ELI5, we employed ROUGE-L scores to assess performance.
@@ 4.2 Baselines
We evaluate publicly available close-sourced LLMs, including GPT-3.5-turbo, GPT-4, and GPT-4o. These models represent state-of-the-art LLMs that are not augmented with external retrieval information. The system prompts and instruction formats are shown in Appendix B.6.
@@
We compare our method against several retrieval-augmented baselines. The baselines are categorized into four groups as follows: RAG-based Baselines: This group includes approaches such as IRCoT Trivedi et al. (2023), ReAct Yao et al. (2023), FLARE Jiang et al. (2023), RA-DIT Lin et al. (2024) and Reward-RAG Nguyen et al. (2024), which leverage agent-like strategies for retrieval-augmented generation. LLaMA2-7B-based Baselines: This category consists of standard retrieval-augmented baselines such as Vanilla-RAG, as well as its enhanced versions with additional components like reranker, SFT, Self-RAG Asai et al. (2023), DRAGIN Su et al. (2024) and Smart-RAG Gao et al. (2024). LLaMA2-13B-based Baselines: Similar to the LLaMA2-7B group, this set includes Vanilla-RAG, Reranker, SFT, and Self-RAG Asai et al. (2023), providing a larger-scale comparison using the LLaMA2-13B model. LLaMA3-8B-based Baselines: In this group, we consider models based on LLaMA3-8B, including Vanilla-RAG and its variations with Reranker and SFT. Additionally, we compare our models with more advanced retrieval-augmented methods such as Auto-RAG Yu et al. (2024a), ChatQA-1.5 Liu et al. (2024b), and RankRAG Yu et al. (2024b).

### v2

@@ Abstract
Retrieval-augmented generation (RAG) systems combine large language models (LLMs) with external knowledge retrieval, making them highly effective for knowledge-intensive tasks. A crucial but often under-explored component of these systems is the reranker. Since irrelevant documents in RAG systems can mislead the generator, the reranker plays a vital role in refining retrieved documents to enhance generation quality and explainability. However, it is challenging to determine the appropriate number of documents ($k$) that the reranker should select: too few may result in missing critical information, while too many introduce noise and inefficiencies. Although recent studies have explored LLM-based rerankers, they primarily leverage internal model knowledge and overlook the rich supervisory signals that LLMs can provide, such as using response quality as feedback for optimizing reranking decisions. In this paper, we propose DynamicRAG, a novel RAG framework where the reranker dynamically adjusts both the order and number of retrieved documents based on the query. We model the reranker as an agent optimized through reinforcement learning (RL), using rewards derived from LLM output quality. Across seven knowledge-intensive datasets, DynamicRAG demonstrates superior performance, achieving state-of-the-art results among models of same parameter sizes. The model, data and code are available at https://github.com/GasolSun36/DynamicRAG.
@@ 1 Introduction
Retrieval-augmented generation (RAG) systems have emerged as a powerful approach for combining the strengths of large language models (LLMs) with external knowledge retrieval. This integration has proven highly effective for addressing knowledge-intensive tasks and incorporating up-to-date information into LLMs, leading to notable performance improvements [17, 24, 14]. RAG systems often suffer from two critical challenges: misleading irrelevant retrieved documents that can distort the generation process, and the ’lost-in-the-middle’ phenomenon where important information gets buried within long lists of retrieved candidates. A crucial, yet often underappreciated, component that addresses these issues is the reranker, which assesses the relevance of retrieved documents. The reranker is critical for improving the quality of generated text and enhancing explainability, thereby serving as an indispensable part of the RAG framework [33, 42]
@@
In RAG systems, the reranker’s primary role is to refine the Top-$N$ documents retrieved by the retriever, selecting the $k$ most relevant ones to enhance the answer quality. However, determining the optimal k remains a challenging problem, as highlighted in previous studies [36, 60]. A $k$ that is too small risks omitting critical information, leading to degraded generation quality, while a larger $k$ may introduce irrelevant content, increasing noise and potentially misleading the generator. Furthermore, incorporating excessively long contexts can reduce both efficiency and effectiveness, further complicating the balance between relevance and performance in RAG systems. Striking the right balance requires adaptive strategies that can dynamically adjust $k$ based on query complexity and document diversity, as shown in Figure 1.
@@
Figure 1: Illustration of our dynamic reranker framework. (a) It represents a RAG system without a reranker, where the system primarily focuses on training LLMs. (b) It represents a RAG system with a reranker, where the reranker is trained to filter the Top-$N$ documents to a fixed Top-$K$, which remains constant for all queries. (c) In contrast, it represents our dynamic reranker, where both the reranker and the generator are trained simultaneously. The dynamic reranker adapts to the difficulty of each query by dynamically determining the value of $k$.
Recent work has demonstrated the effectiveness of LLM-based rerankers [9, 51, 26, 23, 30], which leverage large language models’ capabilities to assess document relevance and improve ranking quality. Other works have also used the understanding capabilities of LLMs through sliding window mechanisms to achieve optimal re-ranking results [51, 5]. While these studies demonstrate the effectiveness of LLM-based rerankers, they typically rely on fixed document selection thresholds ($k$) and fail to dynamically adapt to varying query complexity and retrieval quality. Existing approaches primarily exploit LLMs’ internal knowledge to score documents independently, overlooking a key insight: the actual generation quality when using different numbers of documents provides direct feedback about the optimal $k$. This natural reward signal enables us to apply reinforcement learning, where the LLM’s generation quality serves as the reward for selecting the right number of documents. This supervisory signal—derived from the downstream generation task itself—offers a more principled approach to document selection than static thresholds or isolated relevance scoring. For instance, the quality of an LLM’s response—given a query and a ranked set of documents—serves as a direct indicator of document relevance.
@@
We comprehensively evaluate DynamicRAG on knowledge-intensive tasks across seven datasets, including general QA (NQ [25], TriviaQA [19]), multi-hop reasoning (HotpotQA [57], 2WikimQA [15]), long-form generation (ASQA [49], ELI5 [10]), and fact verification (FEVER [53]). Additionally, we assess the recall results of DynamicRAG’s reranker on NQ and HotpotQA. Experimental results show that DynamicRAG significantly outperforms existing fine-tuned and prompting-based approaches, achieving state-of-the-art (SOTA) performance among models of same size while requiring substantially less training data.
@@ 2.1.2 Encoding Phase
Once the relevant documents have been retrieved, both the query $\mathbf{q}$ and the documents $\{D_{1},D_{2},\dots,D_{k}\}$ are encoded into fixed-size vectors using a neural encoder such as BERT [8], RoBERTa [31], or other transformer-based models. The goal is to obtain a dense representation of both the query and documents. Let $\mathbf{q}_{\text{enc}}$ and $\mathbf{D}_{i}$ represent the encoded query and the encoding of document $D_{i}$, respectively:
@@ 3 DynamicRAG
The DynamicRAG framework consists of three key components: (1) a frozen retriever, (2) a trainable dynamic reranker, and (3) a trainable generator that is optimized to effectively leverage the reranker’s dynamically selected $k$ documents. The retriever retrieves relevant documents from a large corpus, while the reranker dynamically determines both the order and the number of documents to be passed to the generator to produce an answer. The generator then produces the final output based on the reranker’s selected documents. By iteratively training the reranker and generator, DynamicRAG achieves improvements in the overall efficiency and effectiveness of the RAG system.
@@ 3.1 Dynamic Reranking
Traditional reranking approaches rely on static ranking models that determine the relevance of retrieved documents independently of the generation task. These models typically operate within a fixed-length input framework, where a list comprising $n$ documents serves as the input, and the output is a reordered sequence containing the same $n$ documents. This inherent limitation prevents these static models from dynamically adapting to the specific needs of the generation process, particularly in terms of the number and arrangement of the selected documents. In contrast, DynamicRAG uses feedback from the generator to guide the reranking process, allowing it to dynamically adjust both the order and the number of documents selected. This enables the reranker to optimize the input of the generator, maximizing the likelihood of producing high-quality output (In this paper, we only consider the list-wise ranking).
Figure 2: Illustration of the training paradigm for our method. We treat Dynamic Reranker as an Agent, which interacts with the Environment, generating Top-K docs and receiving rewards to optimize itself.
To formalize this operation, the reranking process can be expressed as a single function that directly outputs a reordered subset of the initially retrieved documents. Given a query $\mathbf{q}$ and a set of retrieved documents $D=\{D_{1},D_{2},\dots,D_{k}\}$, the reranker computes:
@@
where $\hat{D}\subset D$ is the selected and reordered subset of documents, and $\theta_{r}$ represents the parameters of the reranker model. This formulation encapsulates both the scoring and selection processes, producing a subset $\hat{D}$ with dynamically adjusted order and size. This dynamic adjustment enables the model to exhibit greater flexibility in accommodating diverse queries and the specific requirements of the generation task.
@@ 3.2 Training with Interacting
In this section, we outline the training process for the Reranker, where the architectural overview of which is schematically depicted in Figure 2. The training begins with behavior cloning, which equips the Reranker with the basic ability to adjust both the order and the number of selected documents. Building upon this foundational model, we further refine the Reranker’s performance through interactions with the environment. This interaction allows the Reranker to collect feedback from multiple trajectories and enhance its decision-making policy. The complete training procedure is illustrated in Figure 3 and is presented in detail in Algorithm 1.
@@ 3.2.1 Behavioral Cloning with Expert Trajectories
Behavioral cloning [56, 61] is used to supervised fine-tuning the Reranker by mimicking expert trajectories, allowing the model to learn the fundamental actions required for effective ranking. In this stage, the Reranker focuses on learning how to predict the correct intermediate actions $a_{t}$ based on the given task and context. To achieve this, the Reranker is trained on a dataset of expert demonstrations, denoted as $D_{e}$. This enables the model to acquire basic instruction-following capabilities and leverage prior knowledge. The training objective is to maximize the likelihood of the expert’s document selection decisions:
$\mathcal{J}_{BC}(\theta)=\mathbb{E}_{(q,D,k^{*})\sim\mathcal{D}_{e}}\left[\log\pi_{\theta}(k|q,D)\right],$ (10)
where $q$ denotes the query, $D=\{d_{1},d_{2},...,d_{N}\}$ represents the set of retrieved documents, $k^{*}$ is the expert-demonstrated optimal number of documents, and $\pi_{\theta}(k|q,D)$ represents the conditional probability of selecting $k$ documents from the candidate set $D$ given the query $q$.
@@ 3.2.2 Optimizing the Reranker via Exploration and Feedback
where $\mathbb{R}(\mathcal{G},i,\tau)$ quantifies the response quality. However, optimizing this objective is challenging due to the complexity of the action space and the inherent inefficiencies in sampling. To address this, we adopt the Direct Preference Optimization (DPO) framework [44], which simplifies the reward optimization with pairwise comparisons of trajectories.
@@ 3.2.3 Reward Function Design
  • Semantic Similarity (SS): Assesses the alignment between $\hat{y}$ and $y_{gt}$ using BERTScore [62]:
@@
  • Textual Fluency (TF): Evaluates fluency using ROUGE [28] metrics:
@@
  • LLM-Based Evaluation (LLM-Eval): Uses LLM-based scoring to assess alignment with task requirements, where $\mathcal{P}$ denotes the scoring prompt, detailed in Appendix C.3:
@@
IRCoT [55] N/A - - 17.4 - - - -
ReAct [58] N/A - - 35.1 - - 62.0 -
RA-DIT [29] $\sim$ 1,129k 43.5 72.8/- 36.6 - - 86.9 -
FLARE* [18] N/A - - - 51.0 41.3 - -
Reward-RAG [36] Unknown 42.2 75.6/80.4 - - - 89.8 -
LLaMA2-7B [54] N/A 17.9 -/42.5 16.6 17.9 19.0 30.0 15.6
@@
Self-RAG (LLaMA2-7B) [1] $\sim$ 150k 36.4 -/66.4 - - 30.0 - -
DRAGIN [50] N/A - - 23.2 22.0 - - -
Smart-RAG (LLaMA2-7B) [12] $\sim$ 218k - - - - 26.6 - -
@@
Auto-RAG (LLaMA3-8B-Instruct) [59] Unknown 37.9 60.9/- - - 30.0 - -
ChatQA-1.5 (LLaMA3-8B) [32] $\sim$ 442k 42.4 81.0/87.6 33.4 26.8 - 90.9 -
RankRAG (LLaMA3-8B) [60] $\sim$ 470k 50.6 82.9/89.5 35.3 31.4 - 93.8 -
@@ 4.1 Datasets and Evaluation Metrics
We evaluate DynamicRAG’s performance on comprehensive knowledge-intensive question-answering tasks, spanning over seven datasets and covering different types of challenges: Natural Questions [25], TriviaQA [19], HotpotQA [57], 2WikimQA [15], FEVER [53], ASQA [49], and ELI5 [10]. For the first five datasets (NQ, TriviaQA, HotpotQA, 2WikimQA, and ASQA), we follow prior studies [1, 60] and adopt exact match as the evaluation metric. For TriviaQA and FEVER, we used accuracy, while for ELI5, we employed ROUGE-L scores to assess performance [60, 41].
@@ 4.2 Baselines
We evaluate publicly available close-sourced LLMs, including GPT-3.5-turbo, GPT-4, and GPT-4o. These models represent state-of-the-art LLMs that are not augmented with external retrieval information. The system prompts and instruction formats are shown in Appendix C.3.
@@
We compare our method against several retrieval-augmented baselines. The baselines are categorized into four groups as follows: RAG-based Baselines: This group includes approaches such as IRCoT [55], ReAct [58], FLARE [18], RA-DIT [29] and Reward-RAG [36], which leverage agent-like strategies for retrieval-augmented generation. LLaMA2-7B-based Baselines: This category consists of standard retrieval-augmented baselines such as Vanilla-RAG, as well as its enhanced versions with additional components like reranker, SFT, Self-RAG [1], DRAGIN [50] and Smart-RAG [12]. LLaMA2-13B-based Baselines: Similar to the LLaMA2-7B group, this set includes Vanilla-RAG, Reranker, SFT, and Self-RAG [1], providing a larger-scale comparison using the LLaMA2-13B model. LLaMA3-8B-based Baselines: In this group, we consider models based on LLaMA3-8B, including Vanilla-RAG and its variations with Reranker and SFT. Additionally, we compare our models with more advanced retrieval-augmented methods such as Auto-RAG [59], ChatQA-1.5 [32], and RankRAG [60].